The sales tax rate for the state of Washington was 8.1%.
What is the state sales tax on a $3,800 car in Washington?

Answer by math_tutor2020(3816)


8.1% = 8.1/100 = 0.081

8.1% of 3800 = 0.081*3800 = 307.80 dollars is the final answer.